I am 37 and alex is 15.
write an equation which models how old in years you will be when you are 3 times as old as the youngest person?

Answer by ankor@dixie-net.com(22740)

I am 37 and alex is 15.
write an equation which models how old in years you will be when you are 3 times as old as the youngest person?
:
Let t = time in yrs for this to happen
:
37 - t = 3(15 - t)
37 - t = 45 - 3t
-t + 3t = 45 - 37
2t = 8
t = 4 yrs ago
:
:
Is this true?
4yrs ago you were 33, and Alex was 11